OpenStudy (anonymous):

A block with mass m1 = 8.6 kg is on an incline with an angle θ = 27.0° with respect to the horizontal. For the first question there is no friction between the incline and the block. 1) When there is no friction, what is the magnitude of the acceleration of the block? 2) Now with friction, the acceleration is measured to be only a = 3.03 m/s2. What is the coefficient of kinetic friction between the incline and the block?

OpenStudy (anonymous):

1) Only force is gravity with downward direction. Thus, magnatude is \[|F|=mg\]. 2) Subtract the two to get: \[mg\sin\theta-\mu mg\cos\theta=ma\]. Check that those cos and sin are right geometrically

OpenStudy (anonymous):

so what is the acceleration?

OpenStudy (anonymous):

Isolate the equation for 'a'.